Is (-9,-7) a solution of 3x-7y=15
Answer by stanbon(75887)

Is (-9,-7) a solution of 3x-7y=15
--------
Substitute to see if you get a true statement::
3*-9 -7*-7 = 15
-27 + 49 = 15
----
22 = 15
-----
False
So, (-9,-7) is not a solution.
------------
Cheers,
Stan H.